OpenStudy (anonymous):

A force does 600 joules of work on a 2kg object that starts out at rest. How fast is it moving after the work is done? A force does 600 joules of work on a 2kg object that starts out at rest. How fast is it moving after the work is done? @Mathematics

OpenStudy (anonymous):

W = (mv^2)/2 So, 600 = (2v^2)/2 600 = v^2 v = 600^(1/2) v = 24.5 m/s

OpenStudy (anonymous):

wow thanks !!!!!!!!!!!!!!!!! I over thought it

OpenStudy (anonymous):

2 and 2 cancel out each other

OpenStudy (anonymous):

hahah yeah i just saw that :P

OpenStudy (anonymous):

can you help me with one more? Do you mind?

OpenStudy (anonymous):

Sure. I will try to help with my knowledge. I am still in high school.

OpenStudy (anonymous):

Awesome , thanks .I'll write it now :)

OpenStudy (anonymous):

4. A child, mass 20kg, jumps off a desk and onto his bed. At the peak of his jump, he is 1.2 meters above his bed. He lands on two springs in the mattress that both have a spring constant 400 N/m. How far does he compress the springs

OpenStudy (anonymous):

I got 11.606 and the correct answer is 13.17 idk what I did wrong

OpenStudy (anonymous):

I am srry, i am not getting the answer that you provided.. :/

OpenStudy (anonymous):

:/ for the formula did you get spring potential + gravitational potential = kinetic work?

OpenStudy (anonymous):

btw, did you realize that he landed on 2x springs.

OpenStudy (anonymous):

Does that matter???????

OpenStudy (anonymous):

damn I'm still not getting it

OpenStudy (anonymous):

i am srry bro, i am not getting it either. :(

OpenStudy (anonymous):

Wanna attempt to help with one more please ? :)

OpenStudy (anonymous):

i will try..

OpenStudy (anonymous):

hahaha sorry here it is , . A spring with spring constant k=100 N/m is compressed 1.2 meters and shoots a block, m=1kg, off of a 1.5 meter tall table. How fast is it moving when it hits the ground

OpenStudy (anonymous):

It says to use the quadratic formula in this one , but I don't see how it would use it.

OpenStudy (anonymous):

Wait never mind it doesn't use the quadratic formulla

OpenStudy (anonymous):

ok, a sec..

OpenStudy (anonymous):

ok, 1/2 m v^2 = 1/2kx^2 Solve for v and you get v = 12 m/s for x direction

OpenStudy (anonymous):

hmmm my teacher got 13.17

OpenStudy (anonymous):

Wait for the last problem I asked what answer did you get? I told you the wrong answer

OpenStudy (anonymous):

If you want the velocity in xy direction, Vy = (2(10)(1.5))^.5 = 5.48 m/s Vxy = (5.48^2 + 12^2) ^.5 = 13.19 m/s

OpenStudy (anonymous):

So the teacher is solving for Vxy.

OpenStudy (anonymous):

How did you find the velocity for vy?

OpenStudy (anonymous):

ok i used this formula ; Vf^2 = Vi ^2 + 2ax

OpenStudy (anonymous):

Ohh I remeber that

OpenStudy (anonymous):

I got 0.77 m for the last problem..

OpenStudy (anonymous):

i might be wrong... =/

OpenStudy (anonymous):

what is the answer for the last one?

OpenStudy (anonymous):

ohhh it was 1.05, but yeah I got that also while taking the two springs in account . But when I used one spring I got 1.08. He uses the quadratic formula for that one but I don't know how

OpenStudy (anonymous):

hmm.. ok i am glad you got the answer.

OpenStudy (anonymous):

I'm off by .2 which worries me :P Cause I didn't even use the quadratic formula

OpenStudy (anonymous):

But I can't really see how you're getting Vy using that formula ?

OpenStudy (anonymous):

lol it's totally fine. Ok So you have the height(1.5), the initial velocity is (0), acceleration is gravity (10), and you just solve for the final velocity then, you will get the velocity for y direction.

OpenStudy (anonymous):

But why to the power of .5?

OpenStudy (anonymous):

oh well i can't type square root here. It basically means square root. sqrt(x) = x^(1/2) = x^.5

OpenStudy (anonymous):

ohh I thought so . thanks :) And for finding Vxy did you use the same formula ? and how?

OpenStudy (anonymous):

nope i used Pythagorean theorem.

OpenStudy (anonymous):

Ohhhhhhhhhh that totally makes sense now man. Thanks so much. You've been a tremendous help. Sorry for bugging you

OpenStudy (anonymous):

It's my pleasure. good luck.. ^.^